OpenStudy (anonymous):

Evaluate the indefinite integral: 7/(x(ln(7x))) I've tried so many times but I just can't get it :(

OpenStudy (turingtest):

let u=ln(7x)

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits \frac{7}{x \ln(7x)}\;dx \quad=\quad \int\limits \frac{1}{\ln(7x)}\left(\frac{7}{x}\;dx\right)\]Make the substitution mentioned above. If you arrange the integral like this, maybe the du will jump out at you a lil easier.

zepdrix (zepdrix):

Oh that didn't help.. I shouldn't had left the 7 in there :P my bad.

OpenStudy (anonymous):

I got \[7\ln(\ln(7x))\] but apparently it's wrong...

zepdrix (zepdrix):

Hmm that looks correct.. lemme see what's going on.

OpenStudy (anonymous):

Thanks!! They gave this hint but I'm not sure how it'll help: NOTE: The absolute value of x has to be entered as abs(x).

zepdrix (zepdrix):

Oh oh oh ok. Yah they're being a little fussy. Recall this integral a sec:\[\Large\bf\sf \int\limits \frac{1}{x}\;dx \quad=\quad \ln|x|+c\]

OpenStudy (anonymous):

Ohh I got it! They wanted me to enter it as\[7\ln \left| \ln \left| 7x \right| \right|\] Thanks for your help!!

OpenStudy (turingtest):

stupid absolute value...

zepdrix (zepdrix):

Cool good job \c:/ I think it just be \(\Large\bf\sf 7ln(abs(ln(7x)))\) Just absolute on the outer log. I could be wrong though -_-